OpenStudy (anonymous):

u + 2 = 4 Can someone tell me how to solve this step by step ?

OpenStudy (anonymous):

u = 2

OpenStudy (anonymous):

\[u-2+2=4-2\\u=2\]

OpenStudy (anonymous):

u + 2 = 4 Subtract 2 from both sides -2 -2 ------------- u = 2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Think of the question like this, "What number, plus 2, equals 4?"